OpenStudy (anonymous):

Find the equation of a line that passes through (-4,-1) (2,3).....how do i figure this out?

OpenStudy (anonymous):

first find the slope of the two given coordinates

OpenStudy (anonymous):

how do i find the slope?

OpenStudy (anonymous):

rise / run

OpenStudy (gitface):

Slope can be found using the formula \[m = \frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\] and the eqaution of the line cam be found using the same formula I showed you earlier.

OpenStudy (anonymous):

thank you! my final exam is tomorrow and i dont know how im gonna remember all these things!

OpenStudy (gitface):

Took me until my first year university to remember these. No clue why.

OpenStudy (anonymous):

so it tells me to place the slope (1/3) and one ordered pair into point slope form

OpenStudy (anonymous):

how do i figure out how to do that. I get the y=1/3x partbut i dont get how to use an ordered pair to put it in point slope

OpenStudy (anonymous):

i think i got it actually. the y-y1 thing right?

OpenStudy (gitface):

Yes use that one.

OpenStudy (anonymous):

i got y=1/3x+7/3 but i dont think that is right because its not coming up in my graphing calculator when i try andcheck it

OpenStudy (gitface):

Your slope should come out to 2/3. \[m = \frac{ y _{1}-y _{2} }{ x _{1}-x _{2} } = \frac{ -1-3 }{ -4-2 }= \frac{ -4 }{ -6 }=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

oh. oops. i did it backwards 3-1 over 2--4...2/-6

OpenStudy (anonymous):

oops 6

OpenStudy (anonymous):

thank you

OpenStudy (gitface):

You're welcome.